<h3 class="LC20lb DKV0Md" style="font-size: 20px; font-weight: normal; margin: 0px 0px 3px; padding: 5px 0px 0px; display: inline-block; line-height: 1.3;">Answer in Financial Math for PK NDINI #182650</h3>


1
Expert's answer
2021-05-21T11:16:34-0400

"FV A=A×(\dfrac{(1+\frac{r}{m)}^{m×n}-1}{\frac{r}{m}})"


future value (Fv) = R330 284,00

Interest rate(r)=12.5%

m=12 (monthly compounding)

Time(n )= 4 years


let the size of monthly payment = A,


"\Rightarrow 330 284,00=A×\dfrac{(1+\frac{0.125}{12})^{12×4}-1}{\frac{0.125}{12}}"


"\Rightarrow 33028400=A×(\dfrac{1.0104167^{48}-1}{0.0104167})"


"\Rightarrow 33028400=A×61.87"


"\Rightarrow A=\dfrac{330 284,00}{61.87}"


"A=533849"